# Advances in the Shannon Capacity of Graphs

### v1

@@ 5.1 First approach
$\sqrt[k]{\alpha((\mathsf{G}+\mathsf{H})^{k})}<\Theta(\mathsf{G}+\mathsf{H}),$
@@
[10] Let $\mathsf{G}$ be a universal graph, and let $\mathsf{K}=\mathsf{G}+\operatorname{\mathsf{C}}_{2k+1}$ with $k\geq 2$ be the disjoint union of a universal graph and an odd cycle of length at least $5$. Then, $\Theta(\mathsf{K})$ is not attained at any finite (strong) power of $\mathsf{K}$.
@@
This follows from Theorem 5.1 by showing that $\Theta(\operatorname{\mathsf{C}}_{2k+1})>\alpha(\operatorname{\mathsf{C}}_{2k+1})$. For an odd cycle graph of length $2k+1$, $\alpha(\operatorname{\mathsf{C}}_{2k+1})=k$. By [28, Theorem 7.1], for every $j,k\in\mathbb{N}$ such that $2\leq j\leq k$
$\alpha(\operatorname{\mathsf{C}}_{2j+1}\boxtimes\operatorname{\mathsf{C}}_{2k+1})=jk+\Big\lfloor\frac{j}{2}\Big\rfloor,$
$\Rightarrow$ $\alpha(\operatorname{\mathsf{C}}_{2k+1}^{2})=k^{2}+\Big\lfloor\frac{k}{2}\Big\rfloor>k^{2}=\alpha(\operatorname{\mathsf{C}}_{2k+1})^{2},$
$\Rightarrow$ $\Theta(\operatorname{\mathsf{C}}_{2k+1})\geq\sqrt{\alpha(\operatorname{\mathsf{C}}_{2k+1}^{2})}>\alpha(\operatorname{\mathsf{C}}_{2k+1}).$
@@
This applies in particular to $\mathsf{K}_{1}+\operatorname{\mathsf{C}}_{5}$ since $\mathsf{K}_{1}$ is universal.
@@ 5.2 Second approach
The next approach relies on a lemma in number theory from 1858 by Dedekind. This lemma is provided as follows, whose simple and elegant proof from [39, p. 309] is presented here for completeness.
@@
For all $\ell\in[r]$, $\Theta(\mathsf{G}_{\ell})=\vartheta(\mathsf{G}_{\ell})$ (this follows from the assumptions on $\mathsf{G}_{\ell}$). Thus, by Corollary 3.1
$\Theta(\mathsf{G})=p(\vartheta(\mathsf{G}_{1},\ldots,\vartheta(\mathsf{G}_{r})).$
Furthermore, if $\mathsf{G}_{\ell}$ is a Kneser graph, then $\vartheta(\mathsf{G}_{\ell})=\binom{n_{\ell}-1}{r_{\ell}-1}\in\mathbb{N}$ (by [11, Theorem 13]), and if $\mathsf{G}$ is self-complementary and vertex-transitive or self-complementary and strongly regular then $\vartheta(\mathsf{G}_{\ell})=\sqrt{n_{\ell}}$ (by [34, Theorem 3.26]). So, for every $\ell\neq\ell_{0}$, $\vartheta(\mathsf{G}_{\ell})\in\mathbb{N}$ by assumption. And, by assumption, $\vartheta(\mathsf{G}_{\ell_{0}})\notin\mathbb{N}$. Next, by the assumptions of the theorem, there exists a monomial in $p$, whose degree in $x_{\ell_{0}}$ is odd, and there exists a monomial in $p$ where the variable $x_{\ell_{0}}$ does not occur. Thus, there exists $a,b\in\mathbb{N}$ such that
$p(\vartheta(\mathsf{G}_{1},\ldots,\vartheta(\mathsf{G}_{r}))=a\sqrt{n_{\ell_{0}}}+b.$
Similarly to the concept of proof in Example 5.1, for every $k\in\mathbb{N}$, there exists $c_{k},d_{k}\in\mathbb{N}$ such that
$p(\vartheta(\mathsf{G}_{1},\ldots,\vartheta(\mathsf{G}_{r}))^{k}=c_{k}\sqrt{n_{\ell_{0}}}+d_{k}.$
Since $\sqrt{n_{\ell_{0}}}\notin\mathbb{N}$ (by assumption), $\sqrt{n_{\ell_{0}}}\notin\mathbb{Q}$ by Lemma 5.1, which implies that
$\Theta(\mathsf{G})^{k}\notin\mathbb{N}.$
Hence, for every $k\in\mathbb{N}$,
$\Theta(\mathsf{G})>\sqrt[k]{\alpha(\mathsf{G}^{k})},$
so the capacity of $\mathsf{G}$ is unattainable by any finite strong power of $\mathsf{G}$. ∎
@@
$\mathsf{G}=\mathsf{G}_{1}+\mathsf{G}_{2}+\ldots+\mathsf{G}_{\ell}$
is not attained at any finite power of $\mathsf{G}$. This follows from Corollary 5.2, as the Paley graphs are self-complementary and strongly regular, and the restrictions on $q_{1},\ldots,q_{\ell}$ guaranty that the conditions of Corollary 5.2 hold.
@@ 5.3 Third approach
In the present last approach, we build a family of connected graphs, whose capacity is not attained at any of their finite powers. In particular, we prove that an infinite family of Tadpole graphs is not attained at a finite power.
@@ 6 The Shannon capacity of q-Kneser graphs
Some of the properties of the $q$-Kneser graphs are next presented (see [44, Proposition 3.4]).
@@
By the version of the Erdös-Ko-Rado Theorem for finite vector spaces (see [25, Theorem 9.8.1]), we get
$\alpha(\operatorname{\mathsf{K}}_{q}(n,k))\leq\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-1}\\ {\hskip-8.5359ptk-1\hskip-14.22636pt}\end{array}\Bigr]_{q}.$
Moreover, we can construct a family of $k$-subspaces of $V(n,q)$ containing a fixed $1$-dimensional subspace of $V(n,q)$. This family of subspaces has $\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-1}\\ {\hskip-8.5359ptk-1\hskip-14.22636pt}\end{array}\Bigr]_{q}$ subspaces (see Definition 6.1 and [25, Theorem 9.8.1]), and it is an independent set in $\operatorname{\mathsf{K}}_{q}(n,k)$, thus
$\alpha(\operatorname{\mathsf{K}}_{q}(n,k))\geq\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-1}\\ {\hskip-8.5359ptk-1\hskip-14.22636pt}\end{array}\Bigr]_{q},$
@@
where $\lambda_{\max}$ and $\lambda_{\min}$ are the largest and smallest eigenvalues of the adjacency matrix of $\operatorname{\mathsf{K}}_{q}(n,k)$. By [26, Theorem 2], these eigenvalues are given by
@@
where using (134) with m=nkm=n-k and $r=k$ simplifies the denominator of (120) to
@@
Finally, combining (120) and (139) gives the simplified form
$\vartheta(\operatorname{\mathsf{K}}_{q}(n,k))$ $=\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}\cdot\frac{q^{k}-1}{q^{n}-1}$
$=\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-1}\\ {\hskip-8.5359ptk-1\hskip-14.22636pt}\end{array}\Bigr]_{q}.$
@@ 7 A new inequality for the capacity of graphs
  • The proof of this theorem was introduced by Dorfler and Imrich (1970), and Mckenzie (1971). See section 7.3 in the comprehensive book on graph products [27] (Theorem 7.14).
  • In a paper by Feigenbaum and Schaffer [40], a polynomial-time algorithm was introduced for finding that unique prime factorization (with respect to strong products).
@@
Using (160) and (161) gives
$\Theta(\mathsf{G}+\overline{\mathsf{G}})\geq 2\sqrt{n}.$
Furthermore, let $\mathsf{G}$ be either (1) self-complementary and vertex-transitive or (2) strongly regular or (3) a conference graph, (4) a Latin square graph or (5) any one of the complements of the graphs in (1)–(4). Then, by Theorem 7.2 and [34, Theorems 3.23, 3.26 and 3.28], inequalities (160) and (161) hold with equality, thus inequality (159) holds with equality. ∎
@@
Inequality (159), together with a subset of the sufficient conditions for its equality in Corollary 7.1, were proved by N. Alon in [7, Theorem 2.1]. Our proof is more simple, relying on a different approach, and the original proof in [7] is also presented in Appendix D. In [7, Theorem 2.1], Alon showed that inequality (159) holds with equality if $\mathsf{G}$ is self-complementary and vertex-transitive; some additional alternative sufficient conditions are provided in Corollary 7.1.
@@
$\Theta(m_{1}\mathsf{G}_{1}\boxtimes\ldots\boxtimes m_{\ell}\mathsf{G}_{\ell})\leq\ell^{-\ell}\,\Theta(m_{1}\mathsf{G}_{1}+\ldots+m_{\ell}\mathsf{G}_{\ell})^{\ell}.$
@@
$K(\underline{\alpha})=\{k\in\mathbb{N}:k\alpha_{j}\in\mathbb{N},\forall{j}\in[\ell]\}.$
@@
$\Theta(\mathsf{G}_{1}^{\alpha_{1}k}\boxtimes\ldots\boxtimes\mathsf{G}_{\ell}^{\alpha_{\ell}k})\leq\exp\left(-kH(\underline{\alpha})\right)\,\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})^{k},$ (163)
@@
$H(\underline{\alpha})\triangleq-\sum_{j=1}^{\ell}\alpha_{j}\log{\alpha_{j}}.$
@@
$A=\mathrm{lcm}(\alpha_{1}k,\ldots,\alpha_{\ell}k).$
Next, for every $1\leq j\leq\ell$, define
$n_{j}=\frac{A}{\alpha_{j}k},$ (164)
$a_{j}=\sum_{i=1}^{j}{\alpha_{i}k},$ (165)
and for every $i$, $1\leq i\leq k$, let $j$ be the index that has $a_{j-1}+1\leq i\leq a_{j}$, and define
$m_{i}=n_{j}.$
Now, by Corollary 7.2 (Note that there are $k$ graphs in the capacity)
$\Theta\left(\mathsf{G}_{1}^{\alpha_{1}k}\boxtimes\ldots\boxtimes\mathsf{G}_{\ell}^{\alpha_{\ell}k}\right)$
$\leq(mk)^{-k}\,\Theta\left(m_{1}\mathsf{G}_{1}+\ldots+m_{a_{1}}\mathsf{G}_{1}+m_{a_{1}+1}\mathsf{G}_{2}+\ldots+m_{k}\mathsf{G}_{\ell}\right)^{k}$
$=(mk)^{-k}\,\Theta\left((m_{1}+\ldots+m_{a_{1}})\mathsf{G}_{1}+\ldots+(m_{a_{\ell-1}+1}+\ldots+m_{k})\mathsf{G}_{\ell}\right)^{k}$
$=(mk)^{-k}\,\Theta(n_{1}\alpha_{1}k\mathsf{G}_{1}+\ldots+n_{\ell}\alpha_{\ell}k\mathsf{G}_{\ell})^{k},$
where
@@
$=\frac{A}{k}\left(\prod_{i=1}^{\ell}{\alpha_{i}^{\alpha_{i}k}}\right)^{-\frac{1}{k}}.$
By (164) and (30)
@@
$=A^{k}\,\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})^{k}.$
Finally, we have
@@
$=\exp\left(-kH(\underline{\alpha})\right)\,\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})^{k}.$
@@
In addition, by Theorem 2.18, inequality (7) holds with equality if $\mathsf{G}$ is vertex-transitive or strongly regular. Furthermore, if $\mathsf{G}$ is a $d$-regular graph, then by [31, Proposiotion 1],
@@
which proves (167). Finally, by [31, Proposition 1], if both $\mathsf{G}$ and $\overline{\mathsf{G}}$ are edge-transitive, or if $\mathsf{G}$ is strongly regular, then the right inequality in (7) holds with equality, and if $\mathsf{G}$ is strongly regular, then the left inequality in (7) holds with equality. ∎

### v2

@@ 5.1 First approach
We present here a new and shortened proof, while the proof from [10] is presented in Section D.
@@
$\sqrt[k]{\alpha((\mathsf{G}+\mathsf{H})^{k})}<\Theta(\mathsf{G}+\mathsf{H}),$ (142)
@@
[10] Let $\mathsf{G}$ be a universal graph, and let $\mathsf{K}=\mathsf{G}+\operatorname{\mathsf{C}}_{2k+1}$ with $k\geq 2$ be the disjoint union of a universal graph and an odd cycle of length at least $5$. Then, $\Theta(\mathsf{K})$ is not attained at any finite strong power of $\mathsf{K}$.
@@
This follows from Theorem 5.1 by showing that $\Theta(\operatorname{\mathsf{C}}_{2k+1})>\alpha(\operatorname{\mathsf{C}}_{2k+1})$. For an odd cycle graph of length $2k+1$, $\alpha(\operatorname{\mathsf{C}}_{2k+1})=k$. By [32, Theorem 7.1], for all integers $j$ and $k$ such that $2\leq j\leq k$,
$\alpha(\operatorname{\mathsf{C}}_{2j+1}\boxtimes\operatorname{\mathsf{C}}_{2k+1})=jk+\Big\lfloor\frac{j}{2}\Big\rfloor,$ (143)
which yields
$\Theta(\operatorname{\mathsf{C}}_{2k+1})$ $\geq\sqrt{\alpha(\operatorname{\mathsf{C}}_{2k+1}^{2})}$
$=\sqrt{k^{2}+\bigg\lfloor\frac{k}{2}\bigg\rfloor}$
$>k$
$=\alpha(\operatorname{\mathsf{C}}_{2k+1}).$ (144)
@@
Corollary 5.1 applies in particular to $\mathsf{K}_{1}+\operatorname{\mathsf{C}}_{5}$ since $\mathsf{K}_{1}$ is universal.
@@ 5.2 Second approach
The next approach is based on a classical number-theoretic lemma established by Dedekind in 1858. The lemma is stated below, and for completeness we present the simple and elegant proof given in [43, p. 309].
@@
For all $\ell\in[r]$, $\Theta(\mathsf{G}_{\ell})=\vartheta(\mathsf{G}_{\ell})$ (this follows from the first assumption on $\mathsf{G}_{\ell}$). Thus, by Corollary 3.1
$\Theta(\mathsf{G})=p\bigl(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{r})\bigr).$ (145)
Furthermore, for $\ell\in[r]$, if $\mathsf{G}_{\ell}$ is a Kneser graph, then $\vartheta(\mathsf{G}_{\ell})=\binom{n_{\ell}-1}{r_{\ell}-1}\in\mathbb{N}$ (by [11, Theorem 13]), and if $\mathsf{G}_{\ell}$ is self-complementary and vertex-transitive or self-complementary and strongly regular, then $\vartheta(\mathsf{G}_{\ell})=\sqrt{n_{\ell}}$ (by [39, Theorem 3.26]). By the second assumption on the graphs, it follows that if $\ell\neq\ell_{0}$, then $\vartheta(\mathsf{G}_{\ell})\in\mathbb{N}$, and $\vartheta(\mathsf{G}_{\ell_{0}})\notin\mathbb{N}$. Next, by the assumptions of the polynomial $p\in\mathbb{N}[x_{1},\ldots,x_{r}]$, there exists a monomial in $p$ whose degree in $x_{\ell_{0}}$ is odd, and there exists a monomial in $p$ where the variable $x_{\ell_{0}}$ does not occur. Thus, there exist $a,b\in\mathbb{N}$ such that
$p\bigl(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{r})\bigr)=a\sqrt{n_{\ell_{0}}}+b.$ (146)
Similarly to the concept of proof in Example 5.1, for every $k\in\mathbb{N}$, it follows from (146) that there exist $c_{k},d_{k}\in\mathbb{N}$ such that
$p\bigl(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{r})\bigr)^{k}=c_{k}\sqrt{n_{\ell_{0}}}+d_{k}.$ (147)
Since $\sqrt{n_{\ell_{0}}}\notin\mathbb{N}$ (by assumption), we have $\sqrt{n_{\ell_{0}}}\notin\mathbb{Q}$ by Lemma 5.1. By (145) and (147), it implies that
$\Theta(\mathsf{G})^{k}\notin\mathbb{N},$ (148)
so, for all $k\in\mathbb{N}$,
$\Theta(\mathsf{G})>\sqrt[k]{\alpha(\mathsf{G}^{k})}.$ (149)
This shows that the Shannon capacity of $\mathsf{G}$ is not attained by the independence number of any finite strong power of $\mathsf{G}$. ∎
@@
$\mathsf{G}=\mathsf{G}_{1}+\mathsf{G}_{2}+\ldots+\mathsf{G}_{\ell}$ (151)
is not attained by the independence number of any finite strong power of $\mathsf{G}$. This follows from Corollary 5.2, as the Paley graphs are self-complementary and strongly regular, and the restrictions on $q_{1},\ldots,q_{\ell}$ guaranty that the conditions of Corollary 5.2 hold.
@@ 5.3 Third approach
In (114), one already observes a novel example of a countably infinite family of connected graphs, where the Shannon capacity of every such graph is not achieved by the independence number of any finite strong power. Indeed, this holds since raising the right-hand side of (114) to any power $k\in\mathbb{N}$ does not result in an integer (similarly to (137)). In the present last approach, we build a generalized family of connected graphs, whose capacity is not attained at any of their finite strong powers. In particular, we prove that an extended infinite family of Tadpole graphs is not attained at a finite strong power.
@@ 6 The Shannon capacity of $q$-Kneser graphs
Some of the properties of the $q$-Kneser graphs are next presented (see [52, Proposition 3.4]).
@@
By the version of the Erdös-Ko-Rado Theorem for finite vector spaces (see [29, Theorem 9.8.1]), we get
$\alpha(\operatorname{\mathsf{K}}_{q}(n,k))\leq\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-1}\\ {\hskip-8.5359ptk-1\hskip-14.22636pt}\end{array}\Bigr]_{q}.$ (178)
Moreover, we can construct a family of $k$-subspaces of $V(n,q)$ containing a fixed $1$-dimensional subspace of $V(n,q)$. This family of subspaces has $\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-1}\\ {\hskip-8.5359ptk-1\hskip-14.22636pt}\end{array}\Bigr]_{q}$ subspaces (see Definition 6.1 and [29, Theorem 9.8.1]), and it is an independent set in $\operatorname{\mathsf{K}}_{q}(n,k)$, thus
$\alpha(\operatorname{\mathsf{K}}_{q}(n,k))\geq\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-1}\\ {\hskip-8.5359ptk-1\hskip-14.22636pt}\end{array}\Bigr]_{q},$ (181)
@@
where $\lambda_{\max}$ and $\lambda_{\min}$ are the largest and smallest eigenvalues of the adjacency matrix of $\operatorname{\mathsf{K}}_{q}(n,k)$. By [30, Theorem 2], these eigenvalues are given by
@@
where using (214) with m=nkm=n-k and $r=k$ simplifies the denominator on the right-hand side of (200) to
@@
Combining (200) and (219) finally gives the simplified form
$\vartheta(\operatorname{\mathsf{K}}_{q}(n,k))$ $=\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn}\\ {\hskip-8.5359ptk\hskip-14.22636pt}\end{array}\Bigr]_{q}\cdot\frac{q^{k}-1}{q^{n}-1}$ (230)
$=\Bigl[\hskip-5.69046pt\begin{array}[]{c}{\hskip 5.69046ptn-1}\\ {\hskip-8.5359ptk-1\hskip-14.22636pt}\end{array}\Bigr]_{q},$ (233)
where (233) holds by (214). ∎
@@ 7 A new inequality for the capacity of graphs
The proof of Theorem 7.1 was introduced by Dörfler and Imrich [44], and Mckenzie [45]. See Section 7.3 in the comprehensive book on graph products [31] (Theorem 7.14). In a paper by Feigenbaum and Schäffer [46], a polynomial-time algorithm was introduced for finding that unique prime factorization (with respect to strong products).
@@
Remark 7.1.
An anonymous reviewer brought to our attention that inequality (237) can be related to equation (1) in the preprint [47], for the complementary graph entropy, via the results of [48, Lemma 1] and [6, Equation (11.2)].
@@
Using (246) and (247) gives (245). Furthermore, let $\mathsf{G}$ be either (1) self-complementary and vertex-transitive or (2) strongly regular or (3) a conference graph, (4) a Latin square graph or (5) any one of the complements of the graphs in (1)–(4). Then, by Theorem 7.2 and [39, Theorems 3.23, 3.26 and 3.28], inequalities (246) and (247) hold with equality, thus inequality (245) holds with equality. ∎
@@
Inequality (245), together with a subset of the sufficient conditions for its equality in Corollary 7.1, were proved by N. Alon in [7, Theorem 2.1]. Our proof is more simple, relying on a different approach, and the original proof in [7] is also presented in Appendix E. In [7, Theorem 2.1], Alon showed that inequality (245) holds with equality if $\mathsf{G}$ is self-complementary and vertex-transitive; some additional alternative sufficient conditions are provided in Corollary 7.1.
@@
$\Theta(m_{1}\mathsf{G}_{1}\boxtimes\ldots\boxtimes m_{\ell}\mathsf{G}_{\ell})\leq\ell^{-\ell}\,\Theta(m_{1}\mathsf{G}_{1}+\ldots+m_{\ell}\mathsf{G}_{\ell})^{\ell}.$ (249)
@@
$K(\underline{\alpha})\triangleq\bigl\{k\in\mathbb{N}:\,\alpha_{j}\,k\in\mathbb{N},\;\forall\,j\in[\ell]\bigr\}.$ (250)
@@
$\Theta(\mathsf{G}_{1}^{\alpha_{1}\,k}\boxtimes\ldots\boxtimes\mathsf{G}_{\ell}^{\alpha_{\ell}\,k})\leq\exp\bigl(-k\,H(\underline{\alpha})\bigr)\,\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})^{k},$ (251)
@@
$H(\underline{\alpha})\triangleq-\sum_{j=1}^{\ell}\alpha_{j}\log{\alpha_{j}},\quad\forall\,\underline{\alpha}:\,\alpha_{1},\ldots,\alpha_{\ell}\geq 0,\;\;\sum_{j=1}^{\ell}\alpha_{j}=1.$ (252)
@@
$A\triangleq\mathrm{lcm}(\alpha_{1}k,\ldots,\alpha_{\ell}k),$
and, for every $j\in[\ell]$, define
$n_{j}\triangleq\frac{A}{\alpha_{j}\,k},$ (253)
$a_{j}\triangleq\sum_{i=1}^{j}\alpha_{i}\,k\in\mathbb{N},$ (254)
with $a_{0}\triangleq 0$. For each $i\in[k]$, let $j\in[\ell]$ be the index that satisfies
$a_{j-1}+1\leq i\leq a_{j},$ (255)
and define
$m_{i}\triangleq n_{j},\quad\forall\,i\in[k].$ (256)
By Corollary 7.2 (note that the strong product in the next line involves $k$ graphs)
$\Theta\bigl(\mathsf{G}_{1}^{\alpha_{1}k}\boxtimes\ldots\boxtimes\mathsf{G}_{\ell}^{\alpha_{\ell}k}\bigr)$
$\leq(mk)^{-k}\,\Theta\bigl(m_{1}\mathsf{G}_{1}+\ldots+m_{a_{1}}\mathsf{G}_{1}+\ldots+m_{a_{\ell-1}+1}\mathsf{G}_{\ell}+\ldots+m_{k}\mathsf{G}_{\ell}\bigr)^{k}$
$=(mk)^{-k}\,\Theta\bigl((m_{1}+\ldots+m_{a_{1}})\,\mathsf{G}_{1}+\ldots+(m_{a_{\ell-1}+1}+\ldots+m_{k})\,\mathsf{G}_{\ell}\bigr)^{k}$
$=(mk)^{-k}\,\Theta(\alpha_{1}\,n_{1}\,k\mathsf{G}_{1}+\ldots+\alpha_{\ell}\,n_{\ell}\,k\mathsf{G}_{\ell})^{k},$ (257)
where the last equality holds since, by (253) and (254),
$m_{1}=\ldots=m_{a_{1}}=n_{1},$ (258)
$m_{a_{1}+1}=\ldots=m_{a_{2}}=n_{2},$ (259)
$\vdots$
$m_{a_{\ell-1}+1}=\ldots=m_{k}=n_{\ell},\quad a_{\ell}=k.$ (260)
so, for every $j\in[\ell]$,
$m_{a_{j-1}+1}+\ldots+m_{a_{j}}$ $=(a_{j}-a_{j-1})\,n_{j}$
$=\alpha_{j}\,n_{j}\,k.$ (261)
Calculation of $m$ in the righthand side of (257) gives, by (253) and (258)–(261),
@@
$=\frac{A}{k}\left(\prod_{i=1}^{\ell}{\alpha_{i}^{\alpha_{i}k}}\right)^{-\frac{1}{k}},$ (262)
and also, by (34) and (253),
@@
$=A^{k}\,\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})^{k}.$ (263)
Finally, combining (257), (262) and (263), we get
@@
$=\exp\bigl(-k\,H(\underline{\alpha})\bigr)\,\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})^{k}.$ (264)
@@
In addition, by Theorem 2.19, inequality (7) holds with equality if $\mathsf{G}$ is vertex-transitive or strongly regular. Furthermore, if $\mathsf{G}$ is a $d$-regular graph, then by [36, Proposiotion 1],
@@
which proves (266). Finally, by [36, Proposition 1], if both $\mathsf{G}$ and $\overline{\mathsf{G}}$ are edge-transitive, or if $\mathsf{G}$ is strongly regular, then the right inequality in (7) holds with equality, and if $\mathsf{G}$ is strongly regular, then the left inequality in (7) holds with equality. ∎